OpenStudy (anonymous):

change the expression to a logarithmic expression in base 10 log 4 20

OpenStudy (anonymous):



OpenStudy (anonymous):

@triciaal @Nnesha @Hero

OpenStudy (anonymous):

\[\log_b(x)=\frac{\log_a(x)}{\log_a(b)}\] this is called the "change of base" formula

OpenStudy (anonymous):

so \[\log_4(20)=\frac{\log(20)}{\log(4)}\]

OpenStudy (anonymous):

thank you :)

OpenStudy (anonymous):

your welcome

OpenStudy (triciaal):

no, you are